OpenStudy (anonymous):

14 - 9x = -8(10 + x)

OpenStudy (kropot72):

First you remove the brackets on the right hand side by multiplying 10 by -8 and x by -8. Can you do that?

OpenStudy (anonymous):

-80 and -8x

OpenStudy (kropot72):

Good! So we have: 14 - 9x = -80 - 8x Now collect all the terms in x to the left hand side by adding 8x to both sides: 14 - 9x + 8x = -80 - 8x + 8x Can you simplify the last equation?

OpenStudy (anonymous):

is the answer 94?

OpenStudy (anonymous):

x=94?

OpenStudy (kropot72):

Yes, the answer is x = 94.